log_3 x^2–((log_3)2x)=2
Answer by nerdybill(5404)

Applying log rules:
log_3 x^2–((log_3)2x)=2
log_3 x^2/(2x)=2
x^2/(2x)= 3^2
x^2/(2x)= 9
x/2= 9
x = 18